OpenStudy (anonymous):

Pls help!!A test consists of 60 multiple choice questions, each with five possible answers, only one of which is correct. Compute μX and σX for the distribution. Round to four decimal places if necessary.

OpenStudy (anonymous):

i have mean=12

OpenStudy (kropot72):

The mean of the binomial distribution is given by np where n is the number of trials and p is the probability of success. \[\mu X=np=60\times \frac{1}{5}=?\] The standard deviation is given by \[\sigma X=\sqrt{npq}=\sqrt{(60\times \frac{1}{5}}\times \frac{4}{5})\] where q is the probability of failure = 1 - p

OpenStudy (kropot72):

The radical sign should continue over all the terms for the standard deviation. The variance will be \[60\times \frac{1}{5}\times \frac{4}{5}\]

OpenStudy (anonymous):

so 4/5 is the 4 questions that would be wrong??

OpenStudy (kropot72):

If the answers are chosen at random the probability of choosing a wrong answer is 4/5

OpenStudy (anonymous):

@kropot72 your awesome

OpenStudy (anonymous):

thanks for clearing it up for me !

OpenStudy (kropot72):

You're welcome :)